# Scene123: One Prompt to 3D Scene Generation via Video-Assisted and Consistency-Enhanced MAE

### v1

@@ 5 Conclusions and limitations
Limitations. Our method is both innovative and effective. However, the optimization process remains an area for potential enhancement, as it is currently constrained by our use of implicit neural fields to enforce physical consistency. In our future research, we aim to explore faster, more easily optimized, and sustainable three-dimensional representations of scenes. These will serve as mechanisms to articulate consistency surface constraints, thus accelerating the generation process.
@@ Appendix A Appendix
A.1 Additional Implementation Details
A.1.1 GAN-based Training Strategy Details
@@
For the incorporated discriminator, we adopt a similar architecture, regularization function, and loss weight as the EG3D model [7], with some distinctions. The vanilla structure of a 3D GAN involves a 3D generator that incorporates a super-resolution module, accompanied by a discriminator that accepts both coarse and fine image inputs [7]. Notably, due to the contextual disparities between text-to-3D and 3D GAN applications, we opt to omit the super-resolution component from the 3D GAN architecture. This choice stems from the consistent need to extract mesh or voxel representations from the 3D model within the context of text-to-3D.
A.2 Additional Qualitative Results
We provide additional qualitative results in Fig. 8, including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes, and fairy-tale scenes with realistic renderings and precise depth details.
Figure 8: More style visualization examples generated by Scene123. We provide additional results including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes, and fairy-tale scenes with realistic renderings and precise depth details.
@@ A.3 User Study
For completeness, we follow previous works [86, 93] and conduct a user study by comparing Scene123 with WonderJourney [93], LucidDreamer [12] under 40 image prompts, 20 image prompts for each baseline. For text prompt input, we compare Scene123 with Text2NeRF [95], ProlificDreamer [86], SJC [84] under 60 text prompts, 20 text prompts for each baseline. The volunteers are shown the generated results of our Scene123 and baselines and asked to choose the better one in terms of fidelity, details and vividness. We collect results from 39 volunteers, yielding 3120 pairwise comparisons. The results are shown in Tab. 3. Our method outperforms all of the baselines.
Method WonderJourney LucidDreamer Text2NeRF ProlicDreamer SJC
Prefer baseline 31.12 23.42 46.27 10.25 5.50
Prefer Scene123 (Ours) 68.88 76.58 53.73 89.75 94.50
@@
A.4 Additional analysis
As depicted in Fig. 9, Scene123 doesn’t strictly require a high-quality support set. The novel views generated by our method are generally superior to the views generated by the image-to-video pipeline. This illustrates the adaptability of our approach to cope with such instances of generation failure.
Figure 9: Comparisons between support set generated by image-to-video pipeline and the generated views of Our method.

### v2

@@ 5 Conclusions and limitations
Limitations. Our method is both innovative and effective. However, the optimization process remains an area for potential enhancement, as it is currently constrained by our use of implicit neural fields to enforce physical consistency. In our future research, we aim to explore faster, more easily optimized, and sustainable 3D scene representations. These will serve as mechanisms to articulate consistency surface constraints, thus accelerating the generation process.
@@ Appendix A Appendix
A.1 Implementation Details
We implement our Scene123 with the Pytorch framework (Paszke et al. 2019) and adopt TensoRF (Chen et al. 2022) as the radiance field. To ensure TensoRF can accommodate scene generation over a large view range, we position the camera near the center of the NeRF bounding box and configure it with outward-facing viewpoints. The dimension of the masked codebook is 2048$\times$16. For only text prompt input, we utilize the stable diffusion model in version 2.0 (Rombach et al. 2022) to generated initial image $\mathbf{I_{0}}$ related to the input prompt. Moreover, for depth estimation, we use the boosting monocular depth estimation method (Miangoleh et al. 2021) with pre-trained LeReS model (Yin et al. 2021) to estimate the depth for each view. For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD (Blattmann et al. 2023a) but finetuned for 25 frame generation. During training, we use the same setting as (Chen et al. 2022) for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{t}=0.001$, $\lambda_{dist}=0.001$.
@@ A.1 Implementation Details
GAN-based Training Strategy Details. For the incorporated discriminator, we adopt a similar architecture, regularization function, and loss weight as the EG3D model (Chan et al. 2022), with some distinctions. The vanilla structure of a 3D GAN involves a 3D generator that incorporates a super-resolution module, accompanied by a discriminator that accepts both coarse and fine image inputs (Chan et al. 2022). Notably, due to the contextual disparities between text-to-3D and 3D GAN applications, we opt to omit the super-resolution component from the 3D GAN architecture. This choice stems from the consistent need to extract mesh or voxel representations from the 3D model within the context of text-to-3D.
The camera motion in the SVD model is sometimes limited (Blattmann et al. 2023a), making it unsuitable for directly training the NeRF model. Instead, we use our 3D-Aware Generative Refinement approach, which leverages a discriminator to optimize the process. In this setup, the video diffusion model’s output is treated as real data, while the 3D neural radiance field model’s renderings are treated as fake data. Despite the small range of camera motion in the images generated by SVD, the adversarial training process provides necessary regularization. The discriminator learns to distinguish between real support set images and fake renderings, thereby pushing the neural radiance field model to produce outputs that are indistinguishable from real images. This continuous adversarial interaction compels the generator to improve by learning finer details and more accurate textures, ultimately enhancing the quality of the 3D scene generation. This method ensures that even with limited camera motion, the GAN loss remains effective and contributes to generating high-quality 3D scenes, as shown in Fig. 7 in the manuscript.
Depth alignment details. We use the depth estimation model $f_{e}$ to estimate the depth map $D_{k}^{E}$ for the initial view $\mathbf{I_{0}}$. Note that, unlike the depth map $D_{0}$ of the initial view, $D_{k}^{E}$ cannot be directly taken as the supervision to update the radiance field since it is predicted independently and could conflict with known depth maps such as $D_{k}^{R}$ in the overlapping regions. To solve this issue, we implement depth alignment to align the estimated depth map to the known depth values in the radiance field (Zhang et al. 2024b).
Due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions (Luo et al. 2020), for example, the estimated depth $D_{k}^{E}$ of the inpainted view may be inconsistent with the depth $D_{k}^{R}$ rendered from NeRF since $D_{k}^{R}$ is constrained by previous known views. The inconsistency is manifested in two aspects: scale difference and value difference. For instance, the distance difference of two pixel-aligned spatial points and the depth value of a specific point could be both different in depth maps estimated from different views. The former is the scale difference and the latter is the value difference. In the case of scale difference, we cannot align both points by shift processing because even if we align the depth value of one of the points, the other point is still misaligned. To eliminate the scale and value differences between the overlapping regions of the rendered depth map $D_{k}^{R}$ and the estimated depth map $D_{k}^{E}$ of the novel view, we introduce a two-stage depth alignment strategy. Specifically, we first globally align these two depth maps by compensating for mean scale and value differences. Then we finetune a pre-trained depth alignment network to produce a locally aligned depth map.
To determine the mean scale and value differences, we first randomly select $M$ pixel pairs from the overlapping regions and deduce their 3D positions under depth $D_{k}^{R}$ and $D_{k}^{E}$, denoted as $\left\{(\mathbf{x}_{j}^{R},\mathbf{x}_{j}^{E})\right\}_{j=1}^{M}$. Next, we calculate the average scaling score $s$ and depth offset $\delta$ to approximate the mean scale and value differences:
$s=\frac{1}{M-1}\sum_{j=1}^{M-1}\frac{\|\mathbf{x}_{j}^{R}-\mathbf{x}_{j+1}^{R}\|_{2}}{\|\mathbf{x}_{j}^{E}-\mathbf{x}_{j+1}^{E}\|_{2}},$ (6)
$\delta=\frac{1}{M}\sum_{j=1}^{M}\left(z\left(\mathbf{x}_{j}^{R}\right)-z\left(\mathbf{\hat{x}}_{j}^{E}\right)\right),$ (7)
where $\mathbf{\hat{x}}_{j}^{E}=s\cdot\mathbf{x}_{j}^{E}$ indicates the scaled point and $z(\mathbf{x})$ represents the depth value of point $\mathbf{x}$. Then $D_{k}^{E}$ can be globally aligned with $D_{k}^{R}$ by $D_{k}^{global}=s\cdot D_{k}^{E}+\delta$.
Since depth maps used in our pipeline are predicted by a network, the differences between $D_{k}^{R}$ and $D_{k}^{E}$ are not linear, that is why the global depth aligning process cannot solve the misalignment problem. To further mitigate the local difference between $D_{k}^{global}$ and $D_{k}^{R}$, we train a pixel-to-pixel network $f_{\psi}$ for nonlinear depth alignment. During optimization of each view, we optimize the parameter $\psi$ of the pre-trained depth alignment network $f_{\psi}$ by minimizing their least square error in the overlapping regions:
$\mathop{\min}_{\psi}\left\|\left(f_{\psi}(D_{k}^{global})-D_{k}^{R}\right)\odot M_{k}\right\|_{2}.$ (8)
Finally, we can derive the locally aligned depth using the optimized depth alignment network: $\hat{D}_{k}=f_{\psi}(D_{k}^{global})$.
Figure 8: Comparisons of SD inpainting and our Consistency-Enhanced MAE in handling detailed complementation.
A.2 Additional analysis
Comparison between SD inpainting and Our Consistency-Enhanced MAE in detail completion. Our Consistency-Enhanced MAE ensures consistency across views by using a learnable codebook to distribute global information to every invisible area. Traditional VAE methods quantize the latent space, leading to artifacts at the boundaries between quantized latents. In contrast, the cross attention mechnism allows for dynamic and flexible feature weighting, better captures global context, and preserves detailed information without the loss associated with quantization. The learnable codebook continuously adapts during training, enhancing the model’s stability and performance in generating high-quality 3D scenes.
However, the stable diffusion (SD) model, which is originally utilized to inpaint images in LucidDreamer (Chung et al. 2023), is less effective for the multi-view complementation task. While SD can achieve complementation through conditional diffusion, it struggles with complex structures and view-consistency tasks, as it is not tailored for this purpose. Although SD performs well in generating new images, it may not be as effective in complementing local details compared to MAE, which is specifically designed for this task. This superiority of our Consistency-Enhanced MAE in handling detailed complementation is evident in Fig. 8.
Video Generation Model promotes our Scene123’s quality. The 3D scenes generated through our 3D-Aware Generative Refinement module are of higher quality than the image-to-video generation pipeline. As depicted in Fig. 9, our Scene123 does not strictly require a high-quality support set. The novel views generated by our method are generally superior to those produced by the image-to-video pipeline. This is because our 3D-Aware Generative Refinement uses a discriminator to manage optimization, distinguishing between real and fake images. This adversarial interaction improves the generator’s ability to learn finer details and more accurate textures, enhancing the overall quality of 3D scene generation. This demonstrates our approach’s adaptability in handling image-to-video generation failures.
@@ A.3 User Study
For completeness, we follow previous works (Wang et al. 2024; Yu et al. 2023) and conduct a user study by comparing Scene123 with WonderJourney (Yu et al. 2023), LucidDreamer (Chung et al. 2023) under 40 image prompts, 20 image prompts for each baseline. For text prompt input, we also compare Scene123 with WonderJourney (Yu et al. 2023), LucidDreamer (Chung et al. 2023) and Text2NeRF (Zhang et al. 2024b) under 60 text prompts, 20 text prompts for each baseline. The participants are shown the generated results of our Scene123 and baselines and asked to choose the better one in terms of fidelity, details and vividness. We collect results from 39 participants, yielding 3120 pairwise comparisons. The results are shown in Tab. 4. Our method outperforms all of the baselines.
Figure 9: Comparisons between support set generated by image-to-video pipeline and the generated views of Our Scene123.
Method WonderJourney LucidDreamer Text2NeRF
Prefer baseline 31.12 23.42 46.27
Prefer Scene123 (Ours) 68.88 76.58 53.73
@@
A.4 More Qualitative Results
We provide more qualitative results in Fig. 10, Fig. 11, Fig. 12, Fig. 13, including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes with realistic renderings and precise depth details.
Part 1 / 4
Figure 10: More qualitative examples generated by our Scene123 from a single image input.
Part 2 / 4
Figure 11: More qualitative examples generated by our Scene123 from a single image input.
Part 3 / 4
Figure 12: More qualitative examples generated by our Scene123 from a text prompt input.
Part 4 / 4
Figure 13: More qualitative examples generated by our Scene123 from a text prompt input.